OpenStudy (anonymous):

An employee earns $17.25 an hour and 10% commission on total sales at the store. Another employee earns $12.95 an hour and 15% commission on total sales at the store. How much in total sales must the store earn in 1 hour for both employees to have an equal amount of pay? @Compassionate @robtobey @Abhisar

OpenStudy (anonymous):

Will medal and fan!

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

dont just give me the answer explain it

OpenStudy (anonymous):

@robtobey you there?

OpenStudy (anonymous):

@robtobey i dont see your comment

OpenStudy (anonymous):

Let t be the total dollar sales in one hour. Solve the following for t: 17.25 + 0.10t = 12.95 + 0.15t t = 86 dollars